Mark and Bert together can finish the job in 4 hours. How fast can Bert finish the job alone if Mark can finish it in 9 hours?
Answer by ptaylor(2198)

Let x=amount of time Bert needs to finish the job working alone
Then Bert works at the rate of 1/x job per hour
And Mark works at the rate of 1/9 job per hour
Mark and Bert together work at the rate of 1/4 job per hour
So, our equation to solve is:
1/x + 1/9 = 1/4 multiply each term by 36x
36+4x=9x subtract 4x from each side
36=5x
x=7 1/5 = 36/5 hours
CK
1/(36/5)+1/9=1/4
5/36 +4/36 =1/4
9/36=1/4
1/4=1/4
Hope this helps--ptaylor